OpenStudy (bbb911):

Find the greatest common factor 8x power of 2, -24x power of 2

OpenStudy (agent0smith):

\[\Large 8x^2 \text{ and } -24x^2\] yes?

OpenStudy (agent0smith):

You could do prime factorization to help you: \[\Large8x^2 = 2*2*2*x*x\] \[\Large -24x^2 = -1*3*2*2*2*x*x\] What factors do they have in common?